OpenStudy (anonymous):

2. A researcher is going to perform a significance test for the difference between two population proportions (H0 : p1 = p2 ) based on the following sample statistics: X1 = 12, n1 = 20, X2 = 15, n2 = 30. Which of the following would be used to compute the value of z for this test?

OpenStudy (anonymous):

From the given info, you can surmise that \(\hat{p}_1=\dfrac{x_1}{n_1}=\dfrac{12}{20}=0.6\) and \(\hat{p}_2=\dfrac{x_2}{n_2}=\dfrac{15}{30}=0.5\). When you're running the test, you'll be using \[Z=\frac{\hat{p}_1-\hat{p}_2}{\sqrt{\frac{\hat{p}_1(1-\hat{p}_1)}{n_1}+\frac{\hat{p}_2(1-\hat{p}_2)}{n_2}}}\]

OpenStudy (anonymous):

Which of these would be more correct? @SithsAndGiggles

OpenStudy (anonymous):

Can you see the attachment?

OpenStudy (anonymous):

Nope, no attachment.

OpenStudy (anonymous):

How about now?

OpenStudy (anonymous):

Thank you for the help.

OpenStudy (anonymous):

Yep. It looks like the question is asking for what standard error you'd be using. The standard error is the denominator of what I typed. The closest one would be C, but the option uses \(x_1\) and \(x_2\) in place of \(n_1\) and \(n_2\), which makes me think answer might very well be E.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

I believe C. I have seen other example where the n-values would be plenty big. Not that you will probably see this answer anyway, haha.